$480,000 is invested at 6%. How much additional amount of money must be invested at 7% so that the total annual income will be $53,300?
Answer by richwmiller(17219)

0.06*480000+0.07*y=53300
28800+0.07*y=53300
0.07*y=24500
y=24500/0.07
y=350000
check
0.06*480000+0.07*350000=53300
ok